OpenStudy (anonymous):

use any problem solving strategy to solve the following problem. a rectangular room has a length that is 4 feet greater than the width. write a polynomial to describe the area of the room

OpenStudy (whpalmer4):

let the length be \(l\) the width can be \(w\) we know that the length is 4 feet greater than the width, so \(l = w+4\) the area of a rectangle is \[A = l*w\] Your polynomial is \[A = l*w = \](fill in the expression we found for \(l\) in place of \(l\))